# Framing Unionization on Facebook: Communication around Representation Elections in the United States

### v1

@@ Abstract
Digital media have become central to how labor unions communicate, organize, and sustain collective action. Yet little is known about how unions’ online discourse relates to concrete outcomes such as representation elections. This study addresses the gap by combining National Labor Relations Board (NLRB) election data with 158k Facebook posts published by U.S. labor unions between 2015 and 2024. We focused on five discourse frames widely recognized in labor and social movement communication research: diagnostic (identifying problems), prognostic (proposing solutions), motivational (mobilizing action), community (emphasizing solidarity), and engagement (promoting interaction). Using a fine-tuned RoBERTa classifier, we systematically annotated unions’ posts and analyzed patterns of frame usage around election events. Our findings showed that diagnostic and community frames dominated union communication overall, but that frame usage varied substantially across organizations. In election cases that unions won, communication leading up to the vote showed an increased use of diagnostic, prognostic, and community frames, followed by a reduction in prognostic and motivational framing after the event—patterns consistent with strategic preparation. By contrast, in lost election cases unions showed little adjustment in their communication, suggesting an absence of tailored communication strategies. By examining variation in message-level framing, the study highlights how communication strategies adapt to organizational contexts, contributing open tools and data and complementing prior research in understanding digital communication of unions and social movements.
@@ Introduction
Digital media has transformed the landscape of collective action. Civil Society Organizations (CSOs) and social movements increasingly rely on online platforms to mobilize participants, coordinate activities, and frame issues to reach broader audiences. Scholars have highlighted the promises of these digital technologies: lowering barriers to participation, enabling new forms of connective action, and amplifying marginalized voices. At the same time, they have cautioned about their limitations, including resource constraints, fragmented attention, and uncertain effects on long-term organizational capacity (Bennett and Segerberg 2012; Earl and Kimport 2011; Tufekci 2017).
@@
Labor unions provide a distinctive case within the ecosystem of CSOs and social movements. As one of the most institutionalized forms of collective organization, they have experienced declining membership and political influence in the United States since the mid-twentieth century. This led scholars to ask whether their revitalization requires not only traditional collective bargaining but also new forms of communication and public engagement (Milkman 2013; Hurd 1998; Voss and Sherman 2000). Social media platforms offer this type of strategic opportunity, especially in today’s fragmented workforce with precarious employment and immigrant labor (Milkman 2020). Unions have indeed turned to digital media as a strategic opportunity: a tool to communicate with members, reach new audiences, and sustain organizational relevance (Lee 1996). Among digital platforms, Facebook has emerged as particularly important (Carneiro and Costa 2022; Ford and Sinpeng 2022). Like other social media, it supports both outward and inward communication, and attracts users whose demographic profiles resemble many unionized workforces. These features make it strategically relevant for labor mobilization (Blanc 2022). Despite this promise, the platform also introduces new complexities, such as uncertainty about which engagement strategies are most effective (Greene and Kirton 2003), and little is known about how union communication on the platform relates to concrete unionization outcomes at scale, particularly in representation elections.
@@
In this study, we combined NLRB election data with a large corpus of Facebook posts by labor unions in the United States from 2015 to 2024 to examine the relationship between online communication and election outcomes. Using natural language processing and computational methods, we analyzed how unions framed their digital communication around representation elections and whether these patterns varied by electoral outcomes [^1] Data and code are available at
@@
RQ2. Does the use of such frames leading up to an election differ for won and lost cases?
@@
Our contributions are threefold: (i) we provide guidelines for annotating discourse frames in unions’ social media text, with possible application to organizations beyond the union context and to platforms beyond Facebook; (ii) we release a fine-tuned, ready-to-use classification model for detecting frames in unions’ communication; and (iii) we advance research at the intersection of framing and labor movement, offering new insights into the relationship between digital communication and unionization dynamics. While much labor movement scholarship often treats unions at the collective level  (Snow, Benford et al. 1988; Benford and Snow 2000; Kelly 1998; Voss and Sherman 2000), we show how digital platforms enable the observation of framing strategies at the level of individual messages and connect them to concrete organizational outcomes. This micro-comparative approach suits well the heterogeneity of communication practices across unions, and allows us to quantify the role of framing as a mechanism of adaptation in periods of organizational change.
@@
Building on this perspective, our analysis of union communication on Facebook revealed several key patterns, many of which are in line with prior literature. We found that unions predominantly emphasized problem identification and community-building, though frame use varied across organizations. The mobilization potential, embodied by a discourse centered on call-to-action messages, remained largely untapped. These findings align with research on union and civil society organizations on social media (Lovejoy and Saxton 2012; Uba and Jansson 2021; Carneiro and Costa 2022). Unions representing broad, multi-industry workforces (industrial unions) tended to emphasize frames for grievances and solutions, whereas unions representing specialized trades (craft unions) relied less on these frames. Prior to elections, in cases that unions ultimately won, communication intensified around problem identification, solution-oriented messaging, and community building, while solution-oriented and call-to-action frames generally declined after the election. In cases that unions lost, by contrast, framing patterns remained relatively stable from pre- to post-election periods, suggesting that strategic adjustments in communication may play a role in shaping election outcomes. This link between communication strategies and organizational outcomes, as well as their adjustment after victory, is also reflected in prior work on union revitalization (Turner 2005) and theories of organizational life cycles of social movements (Christiansen 2009; Tarrow 2022).
@@
Our findings provide actionable insights for labor unions, civil society organizations, and social movements on their digital communication strategies. In particular, using informal platforms such as social media to discuss issues and propose solutions may be more effective than relying on these channels for superficial engagement, which often dominates professionalized social media management. Beyond these practical implications, our results contribute to theoretical debates on framing and collective action by showing how message-level strategies vary across organizations and electoral contexts. By shifting focus from collective-level analyses of movements to the micro-dynamics of discourse on digital platforms, this study highlights how framing might operate as a mechanism of adaptation in periods of organizational change.
@@
Labor unions, like other civil society organizations and social movements, rely on communication to achieve core organizational goals. Nowadays much of this activity occurs in digital spaces, making it critical to understand the communicative strategies unions employ online.
The theory of collective action frames (Snow, Benford et al. 1988; Benford and Snow 2000) offers a central lens for analyzing these strategies. It identifies the following “schemata of interpretation” (Goffman 1974), or functions of social movements’ communication: diagnostic frames, which highlight problems and assign responsibility; prognostic frames, which propose solutions; and motivational frames, which seek to mobilize audiences. These three core frames also parallel the conditions for union revitalization outlined by Kelly (1998): workers must recognize grievances (diagnostic), perceive them as actionable (prognostic), and see unions as legitimate vehicles for change (motivational).
To capture the specificity of union communication online, we extended this framework with two additional categories. First, community frames that emphasize solidarity, shared identity, and collective strength—aspects particularly salient in union contexts (Carneiro and Costa 2022) and consistent with the community cultivation function in nonprofit communication (Lovejoy and Saxton 2012). Second, engagement frames that capture efforts to stimulate interaction, circulation, and participation, reflecting the affordances and dynamics of social media (Muntinga, Moorman, and Smit 2011).
@@
  • Motivational: encouraging participation through calls to action.
  • Community: fostering solidarity and collective identity.
  • Engagement: promoting interaction and audience involvement.
This framework maps prior research onto our operationalization, providing a theoretically grounded basis for analyzing union social media communication.
@@
Our focus is on the central stage of the unionization process: representation elections and their outcomes (win or loss). We collected data on election results from publicly available records of the National Labour Relations Board (NLRB) (National Labor Relations Board 2025), covering all cases for which a representation election (RC cases in the NLRB data) was petitioned and held between January 2015 and December 2024 for labor unions in the United States. From the available variables, we focused on the numeric identifier of each election case, the election date, the names of participating unions, and the union that won the election.
The dataset comprises 11,416 election cases, of which 10,975 involve a single union, 417 involve two unions, and 24 involve three unions competing in the same election. Since much of the information was entered manually in the NLRB data, union names exhibit considerable variation in formatting and granularity. To facilitate analysis of unions’ Facebook (FB) activity, we standardized these names at the national or international level. For example, “Communications Workers of America Local 7250” was mapped to “Communications Workers of America.” To identify a set of national and international unions, we began with the official list of unions affiliated with the American Federation of Labor and Congress of Industrial Organizations (AFL-CIO) [^2] AFL-CIO affiliated unions. https://aflcio.org/about-us/our-unions-and-allies/our-affiliated-unions. Accessed March 3, 2025., consolidating subunits under their parent organizations (e.g., TNG-CWA under CWA) using regular expressions and rule-based assignments. This process yielded 47 unions. For unions not included in the AFL-CIO list, we identified common patterns and acronyms using a combination of regular expressions (e.g., extracting acronyms in parentheses) and manual grouping, resulting in additional 510 potential national and international unions.
To link these organizations with their digital presence, we queried the Bing Search API [^3] Microsoft Bing Search API. https://api.bing.microsoft.com/v7.0/search. Accessed March 3, 2025 with each union name alongside the string “union Facebook” to identify official FB pages. We parsed the results using regular expressions and manually verified their accuracy. This resulted in 66 unions with official FB accounts, of which 40 provided data through the Meta Research API Content Library (Meta Platforms, Inc. 2024). Table D1 lists the unions involved in the study, together with their acronyms and type of structure. In total, we collected 158,238 posts (157,883 of which had text available) published between January 2015 and December 2024, matching the time span of the election data.
@@
We trained a classifier to detect five discourse frames in union Facebook posts: diagnostic, prognostic, motivational, community, and engagement. Aiming at capturing the nuanced usage of such frames in online communication, we set the task to be multi-label. Following common practices in Natural Language Processing (NLP), we first fine-tuned a RoBERTa-base model (Liu et al. 2019) on the text of the 157,883 posts to better capture typical linguistic patterns of labor union communication on Facebook. After collecting annotations for the presence of discourse frames [^4] We also annotated posts for an additional political endorsement label to capture explicit support for political actors, but given its low prevalence, it was not included in the main analysis; see Appendix A and B., we further fine-tuned this model to perform the multi-label classification task. Details on model configuration and computational resources are provided in Appendix B.
@@
Using this codebook, the coders annotated a new sample of 200 posts, achieving a Cohen’s $\kappa$ inter-annotator agreement of 0.8 across labels. Given this high level of agreement, the coders then independently annotated disjoint sets of additional randomly sampled posts. In total, this resulted in 1,151 new posts, each annotated by a single coder.
@@
We randomly split the annotated dataset into training (80%) and test (20%) sets. The RoBERTa model pre-adapted to Facebook language was then fine-tuned on the training set for the multi-label classification task. After evaluating its performance on the test set, we applied the final classifier to the entire corpus to generate frame predictions for all posts.
@@
We examined how unions’ usage of discourse frames differed in election cases that were won versus lost in the days preceding the vote, and how these patterns changed afterward.
@@
To account for sparsity in posting frequency, we aggregated our measurements using a rolling window of $n_{days}=5$. Given an active union on Facebook, for each day and each frame, we calculated (i) the number of posts containing that frame and (ii) the proportion of such posts relative to all posts on that day. To focus on activity around elections, for each election case, we defined two non-overlapping windows:
@@
  • Pre-election: mean proportion of posts expressing a given frame from 7 to 3 days before the election.
@@
  • Post-election: mean proportion from 3 to 7 days after the election.
@@
These windows exclude the election day itself—avoiding overlap caused by the rolling aggregation—while capturing both the immediate lead-up and aftermath.
Detrended Frames Usage and its Change
@@
Observed frame usage patterns may be driven by long-term temporal dynamics rather than immediate responses to specific events (e.g., a gradual increase in certain frames over time). To separate event-specific effects from such background trends, we computed detrended scores. For each union, we identified baseline periods of 18 consecutive days during the reference period (January 2015–December 2024) in which the union was not involved in a specific election outcome (win or loss). We assigned a pseudo-event date at the midpoint of each baseline period and applied the same pre- and post-event aggregation procedure as for real elections.
We then computed the detrended usage of each frame as
${U_{b}}^{D}=U_{b}-{U_{b}}^{baseline},{U_{a}}^{D}=U_{a}-{U_{a}}^{baseline}$
@@
where $U_{b}$ and $U_{a}$ denote frame usage before and after the election, and ${U_{b}}^{baseline}$ and ${U_{a}}^{baseline}$ are the corresponding baseline values. Values of ${U_{b}}^{D}$ or ${U_{a}}^{D}$ close to 0 indicate frame usage similar to baseline periods, positive values indicate higher-than-baseline usage, and negative values indicate lower-than-baseline usage. By examining pre-election usage separately for cases that were won and cases that were lost, we described how unions adjusted their communication relative to the baseline. Comparing the two groups directly allowed us to identify differences in their pre-election communication strategies.
We classified changes in frame usage from the pre- to post-election periods into three categories: decrease, stable, and increase. For each union–case instance, we computed the change in detrended frame usage as the offset
$O={U_{a}}^{D}-{U_{b}}^{D}$
To assign union-case instances to categories of change, we examined the combined distribution of $O$ values (across both won and lost cases). We then used the 25th and 75th percentiles of this distribution as thresholds:
@@
$O=\begin{cases}\textbf{decrease},&O<\text{percentile}_{25}\\ \textbf{stable},&\text{percentile}_{25}\leq O\leq\text{percentile}_{75}\\ \textbf{increase},&O>\text{percentile}_{75}\end{cases}$
@@
This procedure provided a consistent, data-driven way to categorize changes in frame usage, enabling systematic comparisons between won and lost election cases and across frames.
@@
To ensure that results generalize well beyond specific design choices, we conducted several robustness checks: (i) we varied the rolling aggregation window size ($n_{days}=5$ vs $n_{days}=3$) (see Appendix C.1); (ii) we balanced the number of won and lost cases within each labor union, repeating analyses across multiple random seeds (see Appendix C.2); and (iii) we down-sampled highly active unions to the 90th percentile of the distribution of representation cases by labor unions, repeating analyses across multiple random seeds (see Appendix C.2).

### v2

@@ Abstract
Digital media have become central to how labor unions communicate, organize, and sustain collective action. Yet little is known about how unions’ online discourse relates to concrete outcomes such as representation elections. This study addresses the gap by combining National Labor Relations Board (NLRB) election data with 158k Facebook posts published by U.S. labor unions between 2015 and 2024. We focused on five discourse frames widely recognized in labor and social movement communication research: diagnostic (identifying problems), prognostic (proposing solutions), motivational (mobilizing action), community (emphasizing solidarity), and engagement (promoting social media interaction). Using a fine-tuned RoBERTa classifier, we systematically annotated unions’ posts and analyzed patterns of frame usage around election events. Our findings showed that diagnostic and community frames dominated union communication overall, but that frame usage varied substantially across organizations. Greater use of diagnostic, prognostic, and community frames prior to an election was associated with higher odds of a successful outcome. After elections, framing patterns diverged depending on results: after wins, the use of prognostic and motivational frames decreased, whereas after losses, the use of prognostic and engagement frames increased. By examining variation in message-level framing, the study highlights how communication strategies correlate with organizational success, contributing open tools and data, and complementing prior research in understanding digital communication of unions and social movements.
@@ Introduction
Digital media has transformed the landscape of collective action. Civil Society Organizations (CSOs), interest groups, and social movements increasingly rely on online platforms to mobilize participants, coordinate activities, and frame issues to reach broader audiences. Scholars have highlighted the promises of these digital technologies: lowering barriers to participation, enabling new forms of connective action, and amplifying marginalized voices. At the same time, they have cautioned about their limitations, including resource constraints, fragmented attention, and uncertain effects on long-term organizational capacity (Bennett and Segerberg 2012; Earl and Kimport 2011; Tufekci 2017).
@@
Labor unions provide a distinctive case within this ecosystem. As one of the most institutionalized forms of collective organization, they have experienced declining membership and political influence in the United States since the mid-twentieth century. This led scholars to ask whether their revitalization requires not only traditional collective bargaining but also new forms of communication and public engagement (Milkman 2013; Hurd 1998; Voss and Sherman 2000). Given today’s fragmented workforce with precarious employment and immigrant labor (Milkman 2020), social media platforms offer a strategic opening for unions. Accordingly, unions have adopted digital media as a communication infrastructure to mobilize members, reach new audience, and maintain organizational relevance (Lee 1996).
Among digital platforms, Facebook has emerged as particularly important (Carneiro and Costa 2022; Ford and Sinpeng 2022). Like other social media, it supports both outward and inward communication, and attracts users whose demographic profiles resemble many unionized workforces. These features make it strategically relevant for labor mobilization (Blanc 2022). In this study, we focus on Facebook for both substantive and practical reasons. Among the 77 unions affiliated with the AFL-CIO [^1] AFL-CIO affiliated unions. https://aflcio.org/about-us/our-unions-and-allies/our-affiliated-unions. Accessed March 3, 2025. (the largest federation of unions in the United States), 63 link to Facebook, 66 to Twitter, and 43 to YouTube. Although Twitter is slightly more prevalent, its data has become increasingly restricted (Murtfeldt et al. 2024). Facebook therefore offers both broad adoption and comparatively accessible data, making it well suited for empirical analysis. Despite its potential, much remains unknown about how unions use the platform effectively: there is uncertainty about which engagement strategies are most effective (Greene and Kirton 2003), and about how union communication on the platform relates to concrete unionization outcomes at scale.
@@
In this study, we combined National Labor Relations Board (NLRB) election data with a large corpus of Facebook posts by labor unions in the United States from 2015 to 2024 to examine the correlation between online communication and representation election outcomes. Representation elections are administered by the NLRB and allow eligible employees to vote on whether a union should be certified as their exclusive bargaining representative. Such elections occur regularly across industries, initiated by employee petitions. Using natural language processing and computational methods, we analyzed how unions framed their digital communication around representation elections and whether these patterns varied by electoral outcomes [^2] Data and code are available at
@@
RQ2. How is the use of different discourse frames prior to an election associated with the likelihood of electoral success?
@@
While our primary analysis focused on discourse frames, prior research on communication suggests that the emotional content can shape mobilization (Feldman and Hart 2016) and that the topics discussed may influence audience response (Wang and Zhou 2021). We therefore complemented our frame-based analysis with emotion detection and topic modeling to provide a more nuanced view of union communication.
@@
Our contributions are threefold: (i) we provide guidelines for annotating discourse frames in unions’ social media text, (ii) we release a fine-tuned, ready-to-use classification model for detecting frames in unions’ communication; and (iii) we advance research at the intersection of framing and labor movement, offering insights into the relationship between digital communication and unionization dynamics. Although much of labor movement scholarship treats unions at the collective level (Snow, Benford et al. 1988; Benford and Snow 2000; Kelly 1998; Voss and Sherman 2000), we show how digital data enables the observation of framing strategies at the level of individual messages and connect them to concrete organizational outcomes. This micro-comparative approach fits well with the heterogeneity of communication practices across unions, and allows us to study how framing varies during periods of organizational change.
@@
Building on this perspective, our analysis of union communication on Facebook revealed several key patterns. We found that unions predominantly emphasized problem identification and community-building, though frame use varied across organizations. The mobilization potential, embodied by a discourse centered on call-to-action messages, remained largely untapped. This pattern echoes prior research showing that unions and CSOs tend to prioritize problem framing and information dissemination on social media, while making more limited use of call-to-action messages (Lovejoy and Saxton 2012; Uba and Jansson 2021; Carneiro and Costa 2022). Unions representing broad, multi-industry workforces (industrial unions) tended to emphasize frames for grievances and solutions, whereas unions representing specialized trades (craft unions) relied less on these frames. By fitting a Bayesian logistic regression model to predict representation election outcomes, we found that a greater focus on problem identification, solution-oriented messaging, and community building prior to the election is associated with higher odds of winning. In cases that were won, solution-oriented and call-to-action frames generally declined after the election. In cases that were lost, by contrast, solution-oriented and interaction-focused frames increased after the election. Overall, these results indicate that election outcomes are associated with distinct patterns of communicative emphasis. These results are based on observational data and do not allow us to make causal claims about the effect of frames on electoral success. The association between communication strategies and organizational outcomes, as well as their adjustment after victory, is also reflected in prior work on union revitalization (Turner 2005) and theories of organizational life cycles of social movements (Christiansen 2009; Tarrow 2022).
@@
Our findings provide actionable insights for labor unions, CSOs, interest groups, and social movements on their digital communication strategies. In particular, successful election outcomes are more frequently associated with social media use emphasizing issue discussion, the proposal of solutions, and solidarity rather than engagement prompts (e.g., encouraging likes or shares) or mobilization cues. Beyond these practical implications, our results contribute to theoretical debates on framing and collective action by showing how message-level strategies vary across organizations and electoral contexts. By shifting focus from collective-level analyses of movements to the micro-dynamics of discourse on digital platforms, this study offers a lens for examining how framing may be involved in processes of adaptation during periods of organizational change.
@@
Labor unions, like other CSOs, interest groups, and social movements, rely on communication to achieve core organizational goals and lobbying. Nowadays, much of this activity occurs in digital spaces, making it critical to understand the communicative strategies unions employ online.
Framing is central to communication studies, but its meaning varies across traditions. In rhetorical approaches, frames selectively emphasize certain considerations to shape audience interpretations (Gamson and Modigliani 1989; Tversky and Kahneman 1981; Druckman 2001; Kinder 1998). Social movement research treats frames more broadly as recurring points of emphasis that structure problems, propose solutions, and motivate action (Snow, Benford et al. 1988; Benford and Snow 2000), highlighting meaning-making and collective behavior rather than persuasion. Interest group research similarly sees framing as strategic and instrumental (Baumgartner et al. 2009), but emphasizes deployment under specific political or organizational conditions rather than exhaustive typologies. Building on these perspectives, we adopt a functional view of frames in union social media, analyzing a selective subset that reflects their role in organizing meaning and facilitating collective action. Although not exhaustive, these categories provide a theoretically grounded and analytically appropriate framework for studying labor union communication online.
The theory of collective action frames (Snow, Benford et al. 1988; Benford and Snow 2000) offers a central lens for analyzing frames and functions of unions’ social media communication: diagnostic frames, which highlight problems and assign responsibility; prognostic frames, which propose solutions; and motivational frames, which seek to mobilize audiences. These three core frames also parallel the conditions for union revitalization outlined by Kelly (1998): workers must recognize grievances (diagnostic), perceive them as actionable (prognostic), and see unions as legitimate vehicles for change (motivational).
To capture the specificity of union communication online, we draw on literature on organizational and movement communication on social media (e.g., Lovejoy and Saxton 2012; Carneiro and Costa 2022; Muntinga, Moorman, and Smit 2011), which identifies a comprehensive range of frame types. Guided by iterative analysis of the data, we focus on those categories that are both theoretically meaningful and empirically prevalent. Specifically, we examine two frame types. Community frames emphasize shared identity and solidarity without urging action (Carneiro and Costa 2022; Lovejoy and Saxton 2012), while engagement frames promote platform-specific interaction (e.g., liking, sharing), reflecting the affordances of social media (Muntinga, Moorman, and Smit 2011).
@@
  • Motivational: encouraging participation through calls to action relevant to union goals (e.g., urging followers to join a union or attend protests).
  • Community: fostering solidarity and collective identity, without prompting a specific action.
  • Engagement: leveraging social media affordances (likes, shares, comments) to sustain visibility and algorithmic reach, rather than to mobilize action or express collective identity.
This framework maps prior research on framing and union communication onto an operationalization of key framing functions observable in social media discourse.
@@
Our focus is on the central stage of the unionization process: representation elections and their outcomes (win or loss). We collected data on election results from publicly available records of the NLRB (National Labor Relations Board 2025), covering all cases for which a representation election (RC cases in the NLRB data) was petitioned and held between January 2015 and December 2024 for labor unions in the United States. From the available variables, we focused on the numeric identifier of each election case, the election date, the names of participating unions, and the union that won the election.
The dataset comprises 11,416 election cases, of which 10,975 involve a single union, 417 involve two unions, and 24 involve three unions competing in the same election. On average, there were 20 elections per week between 2015 and 2024. Once elected, a union remained the representative for a given set of workers for an average of 237 days. Since much of the information was entered manually in the NLRB data, union names exhibit considerable variation in formatting and granularity. To facilitate analysis of unions’ Facebook (FB) activity, we standardized these names at the national or international level. For example, “Communications Workers of America Local 7250” was mapped to “Communications Workers of America.” To identify a set of national and international unions, we began with the official list of unions affiliated with the AFL-CIO, consolidating subunits under their parent organizations (e.g., TNG-CWA under CWA) using regular expressions and rule-based assignments. This process yielded 47 unions. For unions not included in the AFL-CIO list, we identified common patterns and acronyms using a combination of regular expressions (e.g., extracting acronyms in parentheses) and manual grouping. This resulted in additional 510 potential national and international unions, including a substantial amount of noise from local chapters which we could not easily reconduct to parent unions.
To link these organizations with their digital presence, we queried the Bing Search API [^3] Microsoft Bing Search API. https://api.bing.microsoft.com/v7.0/search. Accessed March 3, 2025 with each union name alongside the string “union Facebook” to identify official FB pages. We parsed the results using regular expressions and manually verified their accuracy. This resulted in 66 unions with official FB pages, of which 40 provided downloadable data through the Meta Research API Content Library (Meta Platforms, Inc. 2024) [^4] According to the API terms, data are only downloadable for pages with at least 15,000 followers or likes.. While the resulting number of unions included in the dataset may appear limited, these unions collectively represent approximately 46% of all union members in the United States in the 2024 fiscal year [^5] Data from https://www.dol.gov/agencies/olms/public-disclosure-room. Table E1 lists the unions involved in the study, along with their acronyms and type of structure. In total, we collected 158,238 posts (157,883 of which had text available) published between January 2015 and December 2024, matching the time span of the election data.
@@
Control Factors
We collected data on union characteristics and social media activity to account for factors that may influence election outcomes.
Traditional union-level variables were drawn from public records published by the U.S. Department of Labor [^6] https://www.dol.gov/agencies/olms/public-disclosure-room and by the Center for Union facts [^7] https://unionfacts.com/find-your-union/. These include the number of members, assets, disbursements, liabilities, and primary industry (e.g., construction, educational services). From the NLRB election data, we recorded the state in which each election occurred to capture broader state-level institutional and political context. Separately, we coded whether the election took place in a right-to-work jurisdiction, capturing a specific legal dimension of state labor regimes. Finally, following a common distinction in labor movement scholarship (Herrigel 1993; Dubofsky 2000), we annotated unions for their organizational type (craft vs. industrial) through manual review of their official websites.
For social media activity, we drew on the collected Facebook data, recording follower counts and the total number of posts over the analyzed period. From the latter, we computed the average number of posts per week. To reduce multicollinearity, we evaluated Variance Inflation Factors (VIFs) and excluded variables with $\text{VIF}\geq 5$, resulting in the removal of union employees and Facebook followers from the final set of controls.
@@
We trained a classifier to detect five discourse frames in union Facebook posts: diagnostic, prognostic, motivational, community, and engagement. Aiming at capturing the nuanced usage of such frames in online communication, we set the task to be multi-label. Following common practices in Natural Language Processing (NLP), we first fine-tuned a RoBERTa-base model (Liu et al. 2019) on the text of the 157,883 posts to better capture typical linguistic patterns of labor union communication on Facebook. After collecting annotations for the presence of discourse frames [^8] We also annotated posts for an additional political endorsement label to capture explicit support for political actors. Given its low prevalence, it was not included in the main analysis., we further fine-tuned this model to perform the multi-label classification task. We also included a RoBERTa-base model directly fine-tuned on the frame detection task, and a BERT-base model as additional classifiers for comparison. Details on model configuration and computational resources are provided in Appendix B.
@@
Using this codebook, the coders annotated a new, random sample of 200 posts, achieving a Cohen’s $\kappa$ inter-annotator agreement of 0.81 averaged across labels [^9] Per-label agreement: 0.88 (diagnostic), 0.86 (prognostic), 0.86 (motivational), 0.62 (community), and 0.85 (engagement). Given this level of agreement, the coders then independently annotated disjoint sets of additional randomly sampled posts. In total, this resulted in 1,151 new posts, each annotated by a single coder. The annotated data is temporally well distributed: posts from 2016 to 2024 each account for between 5% and 14% of the sample, while 2015 contributes just over 2%.
@@
We randomly split the annotated dataset into training (80%) and test (20%) sets. The RoBERTa model, pre-adapted to Facebook language, was then fine-tuned on the training set for the multi-label classification task. After evaluating its performance on the test set, we applied the final classifier to the entire corpus to generate frame predictions for all posts.
Supplementary Linguistic Measures
In addition to coding frames, we also computed two supplementary linguistic measures to characterize posts in more detail. First, we generated an emotional profile for each post based on Ekman’s six basic emotions (Ekman 1992) plus a neutral category, using a pre-trained DistilRoBERTa classifier101010j-hartmann/emotion-english-distilroberta-base on HuggingFace to produce a probability vector over the seven emotional categories. Second, we applied topic modeling to posts using BERTopic (Grootendorst 2022) with DistilRoBERTa embeddings (Sanh et al. 2019).
@@
We examined how unions’ usage of discourse frames prior to elections can contribute to the prediction of elections’ outcomes, and how frame usage changed after the elections.
@@
To account for sparsity in posting frequency, we aggregated daily frame usage over short, overlapping windows of $n_{days}=5$. Given an active union on Facebook, for each day and each frame, we calculated (i) the number of posts containing that frame and (ii) the proportion of such posts relative to all posts on that day. To focus on activity around elections, for each election case, we defined two non-overlapping aggregated measures:
@@
  • $\text{{pre}}_{fi}$: mean proportion of posts expressing frame $f$ from 7 to 3 days before election $i$.
@@
  • $\text{{post}}_{fi}$: mean proportion of posts expressing frame $f$ from 3 to 7 days after election $i$.
@@
These aggregated measures exclude the election day itself—avoiding overlap caused by the moving aggregation—while capturing both the immediate lead-up and aftermath. The aggregation does not introduce temporal smoothing, as analyses rely on fixed pre- and post-election summaries.
Predicting Election Outcome
To examine whether the use of specific frames was associated with electoral success, we fitted a Bayesian logistic regression model to predict the electoral outcome (win vs. loss). In this framework, regression coefficients are estimated as posterior distributions, providing a full distribution of plausible values for each effect. Weakly informative priors were automatically assigned to all coefficients using the bambi Python package. The model was specified as follows:
$\text{logit}\big(\Pr(\text{election\_win}_{i}=1)\big)=\beta_{0}+\sum_{f\in F}\beta_{f}\text{pre}_{fi}$
$+\sum_{e\in E}\beta_{e}\text{emo}_{ei}+\sum_{c\in C}\beta_{c}X_{ci}+u_{\text{union[i]}}+u_{\text{election-state[i]}}$
where $election\_win_{i}$ is the binary indicator of the outcome for election $i$, and $\beta_{0}$ is the intercept. Predictors fall into four groups: framing variables, other linguistic characteristics, traditional union-level controls, and online presence controls.
Framing variables are captured by $\text{pre}_{fi}$, the proportion of posts prior to election $i$ using frame $f$ (diagnostic, prognostic, motivational, community, or engagement), with coefficients $\beta_{f}$. Other linguistic characteristics are represented by $\text{emo}_{ei}$, the mean predicted probability that posts prior to election $i$ expressed emotion $e$ (anger, disgust, fear, joy, sadness, surprise, or neutral), with coefficients $\beta_{e}$. We do not include topic-level predictors in this analysis because most topics are union-specific (see Results).
Control variables $X_{ci}$ with coefficients $\beta_{c}$ include traditional union-level characteristics and online presence measures. These divided into traditional union level controls and online presence controls. Union-level controls comprise union size, assets, disbursements, liabilities, primary industry, organizational structure (craft vs. industrial), and the presence of a right-to-work law in the election state. Online presence is measured by the average number of weekly posts. All numerical controls are standardized. We also included normally distributed random intercepts for unions, $u_{\text{union[i]}}\sim\mathcal{N}(0,\sigma_{u}^{2})$, and election states, $u_{\text{election-state[i]}}\sim\mathcal{N}(0,\sigma_{u}^{2})$, to account for unobserved heterogeneity.
Detrended Frame Usage: Changes Pre- vs. Post-Election
@@
To examine how unions adjusted their communication around elections, we computed detrended frame usage to remove long-term trends. For each union, baseline periods of 18 consecutive days were identified during January 2015–December 2024 when the union was not involved in a specific election. A pseudo-event date was assigned at the midpoint of each baseline period, and pre- and post-event aggregation was applied like for real elections.
We then computed the detrended usage of frame $f$ as
${\text{pre}_{fi}}^{D}$ $=\text{pre}_{fi}-{\text{pre}_{fi}}^{baseline}$
${\text{post}_{fi}}^{D}$ $=\text{post}_{fi}-{\text{post}_{fi}}^{baseline}$
@@
where $\text{pre}_{fi}$ and $\text{post}_{fi}$ denote usage of frame $f$ before and after the election $i$, and ${\text{pre}_{fi}}^{baseline}$ and ${\text{post}_{fi}}^{baseline}$ are the corresponding baseline values. Values of ${\text{pre}_{fi}}^{D}$ or ${\text{post}_{fi}}^{D}$ close to 0 indicate frame usage similar to baseline periods, positive values indicate higher-than-baseline usage, and negative values indicate lower-than-baseline usage.
The change from pre- to post-election was computed as
$O={\text{pre}_{fi}}^{D}-{\text{post}_{fi}}^{D}$
Union-case instances were classified as decrease, stable, or increase using the 25th and 75th percentiles of the $O$ distribution:
@@
$O=\begin{cases}\textbf{decrease},&O\leq\text{percentile}_{25}\\ \textbf{stable},&\text{percentile}_{25}<O\leq\text{percentile}_{75}\\ \textbf{increase},&O>\text{percentile}_{75}\end{cases}$
@@
This procedure provided a consistent way to categorize changes in frame usage, enabling systematic comparisons between won and lost election cases and across frames.
@@
To ensure that results generalize well beyond specific design choices, we conducted several robustness checks: (i) we varied the moving aggregation window size ($n_{days}=5$ vs $n_{days}=3$) (see Appendix D.3); (ii) we balanced the number of won and lost cases within each labor union, repeating analyses across multiple random seeds (see Appendix D.4); and (iii) we down-sampled representation elections from unions with exceptionally high electoral activity, capping the number of cases per union at the 90th percentile of the distribution, to reduce the influence of a small number of highly active unions. Analyses were repeated across multiple random seeds (see Appendix D.4).